OpenStudy (anonymous):

In Bear Creek Bay in July, high tide is at 1:00 pm. The water level at high tide is 7 feet at high tide and 1 foot at low tide. Assuming the next high tide is exactly 12 hours later and the height of the water can be modeled by a cosine curve, find an equation for Bear Creek Bay's water level in July as a function of time (t).

OpenStudy (anonymous):

f(t) = 6 cos pi over 2 t + 2 f(t) = 3 cos pi over 2 t + 4 f(t) = 6 cos pi over 6 t + 2 f(t) = 3 cos pi over 6 t + 4

OpenStudy (calculusfunctions):

Alright so you need an equation in the form\[f(t)=acoskt +c\]where a is the amplitude, c is the average water level and\[k =\frac{ 2\pi }{ period }\]So what are you actually having trouble with?

OpenStudy (anonymous):

i just dont understand what i have to do

OpenStudy (calculusfunctions):

You have to find a, k, and c. Do you know how to do that?

OpenStudy (calculusfunctions):

t is the time in hours and f(t) is the depth of the water in meters.

OpenStudy (calculusfunctions):

\[a =\frac{ Maximum - minimum }{ 2 }\]and\[c =\frac{ Maximum + minimum }{ 2}\]

OpenStudy (anonymous):

so a= 5.5 and c=6.5

OpenStudy (calculusfunctions):

No

OpenStudy (calculusfunctions):

Your question reads high tide is 7 feet and low tide is 1 feet. Sorry I said meters earlier. I meant feet.

OpenStudy (calculusfunctions):

High tide (maximum vale) and low tide (minimum value)

OpenStudy (anonymous):

so its 7-1/2 and 7+1/2?

OpenStudy (calculusfunctions):

Yes. But more accurately, (7 - 1)/2 and (7 + 1) /2.

OpenStudy (anonymous):

ohh ok and then a=3 and c= 4

OpenStudy (calculusfunctions):

You Got It!

OpenStudy (anonymous):

thank you so much!

OpenStudy (calculusfunctions):

And now k = ?

OpenStudy (anonymous):

um would you put 2pi over 3 and 4 but different equations?

OpenStudy (calculusfunctions):

No. The period is the length of time between two successive (consecutive) high tides or two successive low tides.

OpenStudy (calculusfunctions):

Hence what is the period and then\[k =\frac{ 2\pi }{ period }\]

OpenStudy (calculusfunctions):

Read the question carefully! It states the period!

OpenStudy (anonymous):

the period would be 12?

OpenStudy (calculusfunctions):

Correct! Because the question reads "the next high tide occurs 12 hours later." So then what does k = ?

OpenStudy (anonymous):

i got .523 repeating....did i do something wrong i multiplied 2 times 3.14 and divided it by 12

OpenStudy (calculusfunctions):

No Decimals! What is the exact value of k?

OpenStudy (calculusfunctions):

The exact value of k as a reduced fraction is...?

OpenStudy (anonymous):

pi/6?

OpenStudy (calculusfunctions):

Yes!!!

OpenStudy (calculusfunctions):

Now then what is the final equation?

OpenStudy (anonymous):

it would be D

OpenStudy (calculusfunctions):

Beautiful!!

OpenStudy (anonymous):

thanks for helping me!!! i appreciate it so much :)

OpenStudy (calculusfunctions):

See! I knew you could do it!!! Have more faith in yourself! You're welcome.

OpenStudy (calculusfunctions):

Anything else while I'm still logged in?

OpenStudy (anonymous):

For an angle Θ with the point (−20, −21) on its terminating side, what is the value of cosine? negative 20 over 29 negative 21 over 29 −20 −21

OpenStudy (calculusfunctions):

There 's a bunch of stuff missing that I can't see. Could you write it out properly please?

OpenStudy (anonymous):

-20/29 -21/29 -20 -21 the rest of it is all that they gave me

OpenStudy (calculusfunctions):

There's stuff missing from the question itself that I can't see. I don't care about the answer choices! I don't need them. I need the question lol

OpenStudy (anonymous):

For an angle Θ with the point (−20, −21) on its terminating side, what is the value of cosine?

OpenStudy (anonymous):

uhhhh what the heck...

OpenStudy (calculusfunctions):

Is there perhaps a diagram there that is not translating here?

OpenStudy (calculusfunctions):

Do you think you can type out the problem yourself, instead of simply trying to copy and paste?

OpenStudy (anonymous):

for the angle theta with the point (-20,-21) on its terminating side, what is the value of cosine?

OpenStudy (calculusfunctions):

Sorry, I had to step away for a few seconds.

OpenStudy (anonymous):

its ok

OpenStudy (calculusfunctions):

\[\cos \theta =\frac{ adj. side }{ hypotenuse }\]  Now what is the hypotenuse side?